# Constrained volume-difference percolation model on the square lattice

### v1

@@ Abstract
We study a percolation model with restrictions on the opening of sites on the square lattice. In this model, each site $s\in\mathbb{Z}^{2}$ starts closed and an attempt to open it occurs at time $t=t_{s}$, where $(t_{s})_{s\in\mathbb{Z}^{2}}$ is a sequence of independent random variables uniformly distributed on the interval $[0,1]$. The site will open if the volume difference between the two largest clusters adjacent to it is greater than or equal to a constant $r$ or if it has at most one adjacent cluster. Through numerical analysis, we determine the critical threshold $t_{c}(r)$ for various values of $r$, verifying that $t_{c}(r)$ is non-decreasing in $r$ and that there exists a critical value $r_{c}=5$ beyond which percolation does not occur. Additionally, when $t=1$ and $1\leq r\leq 9$, we estimate the averages of the density of open sites, the number of distinct cluster volumes, and the volume of the largest cluster.
@@ 1 Introduction
In this paper, we propose and study a percolation model with restrictions on the difference between the volumes of the clusters adjacent to each site. We try to open the sites sequentially, the attempt will be successful if the volume difference between the two largest clusters adjacent to the site is greater than or equal to a constant $r$ (contrained) or there is at most one cluster adjacent to it. This model will be called Constrained volume-difference percolation model (CVDPM) and is formally defined below:
Figure 1: Part of the square lattice showing the attempt to open site $s$ at time $t=t_{s}$. All sites marked with the symbols $\bullet$ or $\times$ are open. Sites marked with identical symbols belong to the same cluster. Note that $M_{1}(t_{s},s)=|\mathcal{C}_{w}|=12$ and $M_{2}(t_{s},s)=|\mathcal{C}_{u}|=|\mathcal{C}_{v}|=9$. Therefore, site $s$ will open if and only if $r\leq 3$.
@@
When $\psi(1;r)=0$, percolation does not occur for any value of $t$, and we define $t_{c}(r)~{}=~{}\infty$. Note that if $r=0$, this model reduces to OSPM with parameter $p=t$. Therefore, $t_{c}(0)~{}=~{}p_{c}(\mathbb{Z}^{2})=0.592746050786(3)$ [22] (the ordinary site percolation threshold).
Remark 1
@@
is positive, and its occurrence implies that when we attempt to open site $k$, $t=t_{k}$, we will have $|\mathcal{C}_{k-1}|=|\mathcal{C}_{k+1}|=1$. Therefore, $|M_{1}(t_{k},k)-M_{2}(t_{k},k)|=0$ and site $k$ will not open. Since the events $(A_{5k})_{k\geq 0}$ are independent and $\sum_{k\geq 0}\mathbb{P}(A_{5k})=\infty$, we can apply the Borel-Cantelli Lemma (see [23]) to conclude that the probability of infinitely many of them occurring is $1$. Therefore, percolation does not occur for any value of $t\in[0,1]$.
The higher the value of $r$, it seems more likely that we will have fewer open sites. Therefore, if percolation occurs for $r=r_{1}$ and $r=r_{2}$, with $r_{1}<r_{2}$, we expect $t_{c}(r_{1})\leq t_{c}(r_{2})$. In other words, $t_{c}(r)$ is non-decreasing in $r$. Furthermore, this same argument suggests that there likely exists a positive value of $r$, denoted as $r_{c}$, such that if $r>r_{c}$, percolation will not occur and, therefore, the model will not exhibit a phase transition ($t_{c}(r)=\infty$).
Through numerical simulations of this model, we estimate the critical time for several values of $r$, finding evidence that $t_{c}(r)$ is non-decreasing in $r$ and $r_{c}=5$. Furthermore, we determine other properties of the model when $t=1$ and $1\leq r\leq 9$: the average density of open sites ($\overline{\rho}(r)$), the average number of distinct cluster volumes ($\overline{n}(r)$), and the average volume of the largest cluster ($\overline{M}(r)$).
In the CVDPM, when $r>0$, the probability of a site being open or closed depends on the status of other sites (long-range dependence), unlike the OSPM and OBPM. Consequently, algorithms like the Leath Algorithm [24] or the Invasion Percolation Algorithm [25, 26], which assume the independence of each site being open or closed, cannot be used. These algorithms generate only a single cluster rather than an entire lattice configuration and are often employed to estimate the critical point of percolation models due to their low computational time and space complexities.
@@ 3 Results
The critical times $t_{c}(r)$ were estimated using FSS (4) and are presented in Table 1. For $r>5$, the model does not exhibit a phase transition ($t_{c}(r)=\infty$), hence $r_{c}=5$. This absence of phase transition highlights the significant impact of the constraint parameter $r$ on the percolation behavior of the system.
@@
Table 1: Estimated critical times $t_{c}(r)$ for $1\leq r\leq 9$.
@@
Figure 2: Graphs of $\overline{\rho}(r)$ (left panel) and $\overline{M}(r)$ (right panel) as a function of $r$. The estimates obtained for these parameters are presented in Table 2. The error bars are smaller than the symbols.
In Figure 2 (left panel), the plot of $\overline{\rho}(r)$ as a function of $r$ is presented. As $r$ increases, it becomes less likely for each site to open and, therefore, $\overline{\rho}(r)$ decreases.
The percentage of sites in the largest cluster provides insights into the formation of dominant clusters. For $r=1$, $\overline{M}(r)$ comprises $96.69(2)\%$ of all the sites in the graph. As $r$ increases, $\overline{M}(r)$ decreases drastically, with the largest cluster comprising only $0.02(1)\%$ of the sites for $r=9$. This trend underscores the fragmentation effect caused by higher constraints, leading to smaller and more dispersed clusters. The graph of $\overline{M}(r)$ as a function of $r$ is presented in Figure 2 (right panel). The most significant decrease in $\overline{M}(r)$ occurs between the values $r=5$ and $r=6$. This behavior is explained by the fact that percolation does not occur for $r>r_{c}=5$.
The number of distinct cluster volumes $\overline{n}(r)$ reveals the diversity of cluster formations within the network. The graph of $\overline{n}(r)$ as a function of $r$ is presented in Figure 3. For $r=1$, $\overline{n}(r)$ is relatively low, $3.8(6)$, indicating that a few distinct cluster volumes dominate the network. This occurs because the largest cluster dominates nearly the entire graph, leaving only smaller clusters in addition to it.
@@
Figure 3: Graph of $\overline{n}(r)$ as a function of $r$. The estimates obtained for this parameter are presented in Table 2. The error bars are smaller than the symbols.
@@ 4 Conclusion
In conclusion, we have introduced and analyzed the Constrained volume-difference percolation model on the square lattice. In this model, starting with all sites closed, we attempt to open them sequentially. The attempt will be successful if the volume difference between the two largest clusters adjacent to the site is greater than or equal to a constant $r$ or if there is at most one cluster adjacent to the site. Our numerical simulations revealed that the critical time $t_{c}(r)$ increases with the constraint parameter $r$, for $1\leq r\leq 5$, and identified a critical value $r_{c}=5$, beyond which percolation does not occur. Additionally, the average density of open sites $\rho(r)$ and the average proportion of nodes in the largest cluster $M(r)$ decrease with increasing $r$, while the average number of distinct cluster volumes $n(r)$ initially increases up to $r=6$ and then decreases for higher values of $r$. These results confirm that stricter constraints hinder the formation of large clusters in the system.

### v2

# Constrained volume-difference site percolation model on the square lattice

@@ Abstract
We study a percolation model with restrictions on the opening of sites on the square lattice. In this model, each site $s\in\mathbb{Z}^{2}$ starts closed and an attempt to open it occurs at time $t=t_{s}$, where $(t_{s})_{s\in\mathbb{Z}^{2}}$ is a sequence of independent random variables uniformly distributed on the interval $[0,1]$. The site will open if the volume difference between the two largest clusters adjacent to it is greater than or equal to a constant $r$ or if it has at most one adjacent cluster. Through numerical analysis, we determine the critical threshold $t_{c}(r)$ for various values of $r$, verifying that $t_{c}(r)$ is non-decreasing in $r$ and that there exists a critical value $r_{c}=5$ beyond which percolation does not occur. Additionally, we find that the correlation length exponent of this model is equal to that of the ordinary percolation model. For $t=1$ and $1\leq r\leq 9$, we estimate the averages of the density of open sites, the number of distinct cluster volumes, and the volume of the largest cluster.
@@ 1 Introduction
In this paper, we propose and study a percolation model with restrictions on the difference between the volumes of the clusters adjacent to each site. We try to open the sites sequentially, the attempt will be successful if the volume difference between the two largest clusters adjacent to the site is greater than or equal to a constant $r$ (contrained) or there is at most one cluster adjacent to it. The proposed model can be motivated by processes involving the spread and interaction of information or states within a networked system, where local dynamics are influenced by the relative sizes of interacting groups. For instance, such dynamics may occur in contexts like the diffusion of opinions, innovations, or behaviors, where an individual (site) adopts a state only if the influence of one group sufficiently outweighs that of another.
This model will be called Constrained volume-difference site percolation model and is formally defined below:
Figure 1: Part of the square lattice showing the attempt to open site $s$ at time $t=t_{s}$. All sites marked with the symbols $\bullet$ or $\times$ are open. Sites marked with identical symbols belong to the same cluster. Note that $M_{1}(t_{s},s)=|\mathcal{C}_{w}|=12$ and $M_{2}(t_{s},s)=|\mathcal{C}_{u}|=|\mathcal{C}_{v}|=9$. Therefore, since $|M_{1}(t_{s},s)-M_{2}(t_{s},s)|=3$, the site $s$ will open if and only if the restriction $r$ chosen for the model satisfies $r\leq 3$.
@@
Note that for $r=0$, the condition $|M_{1}(t_{s},s)-M_{2}(t_{s},s)|\geq r$ is always satisfied. In this case, the model simplifies to OSPM, where sites open independently based on their assigned times $t_{s}$. For $r=1$, a site $s$ opens if the largest cluster adjacent to it exceeds the second-largest cluster by at least one site in volume, or if the second-largest cluster is empty ($M_{2}(t_{s},s)=0$). For $r=2$, this difference must be at least two sites, and so on for larger values of $r$.
@@
When $\psi(1;r)=0$, percolation does not occur for any value of $t$, and we define $t_{c}(r)~{}=~{}\infty$. If $r=0$, this model reduces to OSPM with parameter $p=t$. Therefore, $t_{c}(0)=p_{c}(\mathbb{Z}^{2})=0.592746050786(3)$ [22] (the ordinary site percolation threshold).
@@
is positive, $\mathbb{P}(A_{k})=\delta>0$, and its occurrence implies that when we attempt to open site $k$, $t=t_{k}$, we will have $|\mathcal{C}_{k-1}|=|\mathcal{C}_{k+1}|=1$. Therefore, $|M_{1}(t_{k},k)-M_{2}(t_{k},k)|=0$ and site $k$ will not open. Since the event $A_{k}$ depends only on the five times $t_{k-2},t_{k-1},t_{k},t_{k+1},t_{k+2}$, the events $(A_{5k})_{k\geq 0}$ are independent. Furthermore, since $\sum_{k\geq 0}\mathbb{P}(A_{5k})=\sum_{k\geq 0}\delta=\infty$, the second Borel-Cantelli Lemma (see [23]) ensures that infinitely many events $A_{5k}$ occur with probability $1$, thereby preventing percolation for any $t\in[0,1]$.
The higher the value of $r$, it is intuitive to expect that we will have fewer open sites. Therefore, we conjecture that if percolation occurs for $r=r_{1}$ and $r=r_{2}$, with $r_{1}<r_{2}$, then $t_{c}(r_{1})\leq t_{c}(r_{2})$. In other words, $t_{c}(r)$ is non-decreasing with $r$. Furthermore, this argument suggests the existence of a critical value $r_{c}$, such that for $r>r_{c}$, percolation does not occur and, thus, the model does not exhibit a phase transition ($t_{c}(r)=\infty$).
Through numerical simulations of this model, we estimate the critical time for several values of $r$, finding evidence that $t_{c}(r)$ is indeed non-decreasing with $r$ and that $r_{c}=5$. We also found that the correlation length exponent $\nu$ is equal to that of the ordinary percolation model. Furthermore, we determine other properties of the model when $t=1$ and $1\leq r\leq 9$: the average density of open sites ($\overline{\rho}(r)$), the average number of distinct cluster volumes ($\overline{n}(r)$), and the average volume of the largest cluster ($\overline{M}(r)$).
In the Constrained volume-difference site percolation model, when $r>0$, the probability of a site being open or closed depends on the status of other sites (long-range dependence), unlike the OSPM and OBPM. Consequently, algorithms like the Leath Algorithm [24] or the Invasion Percolation Algorithm [25, 26], which assume the independence of each site being open or closed, cannot be used. These algorithms generate only a single cluster rather than an entire lattice configuration and are often employed to estimate the critical point of percolation models due to their low computational time and space complexities.
@@ 2 Numerical Procedure
$\mathrm{Max}\left(\psi^{\prime}_{L}(t;r)\right)\sim L^{\frac{1}{\nu(r)}},$ (5)
where the left-hand side represents the maximum derivative of $\psi_{L}(t;r)$ with respect to $t$. The uncertainties in $t_{c}(r)$ and $\nu(r)$ were determined from the standard deviation of regression residuals.
@@ 3 Results
The values of $\nu(r)$, presented in Table 1, are consistent with the correlation length exponent for ordinary percolation on the square lattice, $\nu=\frac{4}{3}$ [29], providing evidence that both models belong to the same universality class.
@@
$\nu(r)$ 1.34(1) 1.34(1) 1.34(1) 1.34(1) 1.33(1) --
Table 1: Estimated critical times $t_{c}(r)$, for $1\leq r\leq 9$, and $\nu(r)$, for $1\leq r\leq 5$.
The graph obtained through the relation (5) to estimate $\nu(r)$ for $r=1$ is shown in Fig. 2. Graphs for $2\leq r\leq 5$ are omitted as they are qualitatively similar to the case $r=1$.
Figure 2: Graph obtained through a log-log plot using the FSS relation (5) for the case $r=1$. The slope of the linear regression corresponds to the estimated value of $\frac{1}{\nu(r)}$.
The critical times $t_{c}(r)$ were estimated using the FSS relation (4), and the graph for $r=1$ is shown in Fig. 3. In Fig. 4, we present the behavior of $\psi_{L}(t;1)$ for all $L$, where we observe that the curves intersect near the critical value $t_{c}(1)=0.633306(4)$.
Figure 3: Graph obtained through the FSS relation (4) to estimate the critical time $t_{c}(r)$ for the case $r=1$. The linear coefficient of the line obtained from the linear fit is the estimate for $t_{c}(1)$.
Figure 4: Graphs of the $\psi_{L}(t;r)$ functions for $r=1$ and all values of $L$. The dashed line indicates the estimated critical time $t_{c}(r)=0.633306(4)$. The slope of the graph near the critical time increases as $L$ grows.
We assume $\nu(r)=\frac{4}{3}$ for all $r$ and the estimated values of $t_{c}(r)$ are presented in Table 1. For $r>5$, the model does not exhibit a phase transition ($t_{c}(r)=\infty$), hence $r_{c}=5$. This absence of phase transition highlights the significant impact of the constraint parameter $r$ on the percolation behavior of the system.
@@
As $r$ increases, it becomes less likely for each site to open and, therefore, $\overline{\rho}(r)$ decreases. It is worth noting that our model can be interpreted as a Random Sequential Adsorption model (RSA). In such models, particles are sequentially added to a surface according to specific rules, and the probability of a particle adsorbing onto the surface depends on the spatial arrangement of previously adsorbed particles [30]. An important feature of RSA is the jamming limit, which corresponds to the average density of open sites on the graph at the end of the deposition process. In the Constrained volume-difference site percolation model, this limit is precisely given by the value of $\overline{\rho}(r)$.
The percentage of sites in the largest cluster provides insights into the formation of dominant clusters. For $r=1$, $\overline{M}(r)$ comprises $96.69(2)\%$ of all the sites in the graph. As $r$ increases, $\overline{M}(r)$ decreases drastically, with the largest cluster comprising only $0.02(1)\%$ of the sites for $r=9$. This trend underscores the fragmentation effect caused by higher constraints, leading to smaller and more dispersed clusters. The most significant decrease in $\overline{M}(r)$ occurs between the values $r=5$ and $r=6$. This behavior is explained by the fact that percolation does not occur for $r>r_{c}=5$.
The number of distinct cluster volumes $\overline{n}(r)$ reveals the diversity of cluster formations within the network. For $r=1$, $\overline{n}(r)$ is relatively low, $3.8(6)$, indicating that a few distinct cluster volumes dominate the network. This occurs because the largest cluster dominates nearly the entire graph, leaving only smaller clusters in addition to it.
@@ 4 Conclusion
In conclusion, we have introduced and analyzed the Constrained volume-difference site percolation model on the square lattice. In this model, starting with all sites closed, we attempt to open them sequentially. The attempt will be successful if the volume difference between the two largest clusters adjacent to the site is greater than or equal to a constant $r$ or if there is at most one cluster adjacent to the site. Our numerical simulations revealed that the critical time $t_{c}(r)$ increases with the constraint parameter $r$, for $1\leq r\leq 5$, and identified a critical value $r_{c}=5$, beyond which percolation does not occur. We also found that the correlation length exponent $\nu$ of this model is equal to that of the ordinary percolation model.
Additionally, the average density of open sites $\overline{\rho}(r)$ and the average proportion of nodes in the largest cluster $\overline{M}(r)$ decrease with increasing $r$, while the average number of distinct cluster volumes $\overline{n}(r)$ initially increases up to $r=6$ and then decreases for higher values of $r$. These results confirm that stricter constraints hinder the formation of large clusters in the system.